# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

### v1

@@ 2.1 Survey Design
Under (D2)-(D4), we have that w.r.t the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D1)–(D4). Two–stage designs that satisfy a counter-part of (D1)–(D4) at each of the two stages satisfy also (D1)–(D4) at the subject level (Lemma A.2, Supplementary Material).
@@ 2.2 Embedding of Measures
Kernel mean embeddings (KMEs) [38, 85] provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) [1]. Since these representations are defined purely through expectations of the kernel evaluations, all operations on distributions are reduced to linear operations in the RKHS. As a result, questions about probability measures–such as whether two random variables share the same distribution–can be reformulated as questions about distances between vectors and efficiently evaluated via the kernel trick. The distance between two embedded distributions in the RKHS is called the maximum mean discrepancy (MMD). It is a fully nonparametric way to compare distributions, and for characteristic kernels it is equal to zero only when the two distributions are the same. This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., [86, 87, 37, 80, 33, 67].
From the perspective of constructing random forests for distributional targets, incorporating KMEs into the algorithm offers several practical advantages: (i) It provides a distribution-level target, enabling random forests to discriminate conditional laws via an MMD-based splitting criterion. (ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size [32]. (iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data [62]). As a result, modifications to a standard random forest algorithm are minimal, even for outcomes in abstract spaces. (iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to both finite-population and super-population conditional laws under standard regularity conditions, allowing us to analyze a broad class of sampling designs within a single metric framework.
We now introduce the corresponding concepts and definitions from a mathematical perspective.
Let $\mathbf{k}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a bounded positive definite kernel with RKHS $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ and the associated canonical feature map $\Phi_{\mathbf{k}}(y):=\mathbf{k}(y,\cdot)\in\mathcal{H}$. For any $P\in\mathcal{M}_{b}(\mathcal{Y})$ in the space of finite signed Radon measures, define the kernel mean embedding as the mapping
@@
For our consistency arguments, we require, in addition, that $d_{\mathbf{k}}$ metrizes weak convergence on $\mathcal{M}_{b}(\mathcal{Y})$. That is, for any sequence $P_{1},\dots\in\mathcal{M}_{b}(\mathcal{Y})$, $P_{N}\rightsquigarrow P\iff d_{\mathbf{k}}(P_{N},\;P)\to 0$, as $N\to\infty$. We call kernels that satisfy this property universal (initially proposed in [88]). This property concerns the approximation ability of the KME sample estimate to the true target as $N\to\infty$. We place the following assumptions:
@@
    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of [86]).
@@
    $\mathbf{k}$ is $c_{0}$-universal. Meaning that the RKHS $\mathcal{H}$ induced by $\bm{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t. the uniform norm (Definition 2, Theorem 1 in [20]).
@@
Assumptions (S1)–(K2) ensure that $\mu_{\mathbf{k}}(P)$ is well-defined for $P\in\mathcal{M}_{b}(\mathcal{Y})$ and that $d_{\mathbf{k}}$ is both discriminative (characteristic) and topologically compatible with weak convergence (metrizing). Assumption (K3) is not needed for the statistical theory, but when satisfied, admits a simple finite-dimensional approximation of the MMD through Fourier-feature representation for fast computation. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ ([38]); More details are provided in Section A.10 of the Supplementary Material.
Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding; see Section A.10.
@@ 3 SDRF Algorithm and Theoretical Result
Random forests are machine learning methods that combine a CART-style [27] tree ensemble with a bagging step [15], and can also be viewed as kernel-type estimators (see, e.g. [54, 72]). Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems [17], and several works have documented their strong empirical performance [26, 31]. More recently, conditional distributional versions of random forests based on kernel mean embeddings have been proposed for i.i.d. data [21]. Our proposed method, SDRF, constructs a kernel estimator $\widehat{\omega}_{\mathbf{x}}(\cdot)$ that directly targets conditional distributions and appropriately incorporates the survey design. Given the estimated neighborhood and weights (which together define $\widehat{\omega}_{\mathbf{x}}(\cdot)$), we can directly construct an estimator of the conditional distribution $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$, and hence a plug-in estimator for $\Psi(\mathbf{x})$, namely $\Psi(\mathbf{x})$: $\Psi(\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x}))$.
@@
4:  Draw $\forall k(i)$, $I(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
@@
11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
@@ 3.1 Bagging and Honesty
Resampling with Design
Bagging is a key ingredient of random forests [15], but its study is primarily conducted in an i.i.d. setting. In the i.i.d. case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators [16, 18].
Under a survey design, however, Efron’s classical i.i.d. bootstrap [29] can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
We adopt a design–aware bootstrap (BS) method in line with [95] as the default resampling method for SDRF. It generates multipliers $n^{*}$ that determine the additional up– or down– weighting applied to the sampled distribution, and satisfy properties (R1)–(R3) as detailed in Lemma 3.1.
@@
For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section A.5.1 of the Supplementary Material in the spirit of [95] (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
  1. On resample multipliers $n^{*}$
@@
    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains known cluster and strata structure. Each stratum is re–sampled separately, and if the design has two stages, the second stage resampling multipliers are independent across distinct PSUs conditional on first–stage resamples.
@@
The method by [95] belongs to the class of pseudo-population-based resampling methods. To summarize: For each re-sample, we first use design information such as $\{w_{i}\}_{i\in[n_{s}]}$ to re-construct via the multinomial distribution a pseudo finite population, with expected size equal to $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$; then we re-apply the sampling design $\mathbf{p}$ on the re-constructed finite population to obtain re-samples, thereby respecting the design structure. We refer interested readers to [63, 23] for a review of common resampling methods under design.
Honesty with Design
For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ ([3, 93, 4]). With i.i.d. data, any disjoint partition of the observations yields honesty by construction. However, data from design carries possible dependency across subjects, so separation must be made at the PSU level to ensure the maximum level of independence conditioning on $\xi$. This motivates the introduction of $q\in(0,\,1)$. Alternatively, if one has access to data obtained from a continuous survey under the same design and is willing to assume that $P_{Y|X}$ stays unchanged between two consecutive time points, it is also possible to directly use one sample as $D_{est}$ and the other as $D_{split}$.
In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, [49]) decreases as depth increases. As we discuss later, the performance of the MMD–based split criterion hinges on the effective sample size of $D_{split}$. In this case, one might consider spending the PSUs gradually over the tree growing process. See Algorithm 2, Supplementary Material for details.
@@ 3.2 Splitting Rule
The splitting criterion of an RF algorithm guides the formation of tree structure by determining how the feature space should be partitioned into increasingly homogeneous child nodes. Consider any node where a split is possible (has not met the stopping criteria). Denote $P_{a}\subseteq\mathcal{X}$ the feature space of that node. A candidate split can be parametrized as $\theta=(j,\;t)\in\Pi$, $j=1,\dots,p$, $t\in Range(X_{j})\subseteq[a_{j},\;b_{j}]$. This partitions $P_{a}$ into child regions $C_{L}(\theta):=\{x\in P_{a}:x_{j}\leq t\}$ and $C_{R}(\theta):=\{x\in P_{a}:\;x_{j}>t\}$. With RF considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via Hájek–type estimator. We find the split by maximizing the MMD between the KMEs of the distribution estimates between the two child regions. Formally, $\hat{\theta}=\arg\max_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)$, where
$\begin{split}M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},\end{split}$ (1)
with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}I(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$. When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if exists. We first provide a theorem justifying the choice of this split criterion.
@@
Fix a parent region $P_{a}\subset\mathcal{X}$ at depth $d$ and consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which defines child regions $C_{L}(\theta)$ and $C_{R}(\theta)$. Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score in (1), and let $M^{N}(\theta)$ be its finite-population analog. Suppose (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3. Suppose further $\mathcal{X}\subset\mathbb{R}^{p}$ is bounded. Then, conditional on the realized finite population,
@@
If, in addition, the finite population is an i.i.d. realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}$ and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta_{0}$ under the super-population model).
@@
Theorem 3.2 provides a local result: for a fixed parent node $P_{a}$, the empirical MMD criterion $M^{*}_{n_{s}}(\theta)$ uniformly approximates its finite-population oracle counterpart $M^{N}(\theta)$ and also the super-population score $M(\theta)$. The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique ([90]). The condition $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$ makes explicit that random feature subsampling (via mtrymtry) does not distort the split choice asymptotically, provided the candidate set becomes sufficiently dense around $\theta_{0}$ in expectation. Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to that if using the full finite population, which is in particular assumption (B2) of Theorem 3.4 required for consistency in MMD metric.
One might wonder that, since resamples are determined by non–negative multipliers $n^{*}$, if one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample. This can help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees. Proposition 3.3 below provides a detailed characterization of its consequence.

### v2

@@ 2.1 Survey Design
Under (D2)-(D4), we have that with respect to the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D2)–(D4) (Lemma S1, Supplementary Material). Two–stage designs that satisfy a counterpart of (D2)–(D4) at each of the two stages satisfy also (D2)–(D4) at the subject level (Lemma S2, Supplementary Material).
@@ 2.2 Embedding of Measures
Kernel mean embeddings (KMEs) Gretton_2007; Smola_2007 provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) Aronszajn1950. Since these representations are defined purely through expectations of the kernel evaluations, all operations on distributions are reduced to linear operations in the RKHS. As a result, questions about probability measures–such as whether two random variables share the same distribution–can be reformulated as questions about distances between vectors and efficiently evaluated via the kernel trick. The distance between two embedded distributions in the RKHS is called the maximum mean discrepancy (MMD). It is a fully nonparametric way to compare distributions, and for characteristic kernels it is equal to zero only when the two distributions are the same. This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., Sriperumbudur2010-xr; sripernumbudur_hilbert_2010; gretton2012kernel; sejdinovic_equivalence_2013; francca2020kernel; matabuena2025kernel.
From the perspective of constructing random forests for distributional targets, incorporating KMEs into the algorithm offers several practical advantages: (i) It provides a distribution-level target, enabling random forests to discriminate conditional laws via an MMD-based splitting criterion. (ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size Fournier2015-mu. (iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data lyons2020strong). As a result, modifications to a standard random forest algorithm are minimal, even for outcomes in abstract spaces. (iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to finite-population and super-population conditional laws under standard regularity conditions, allowing for the analysis of a broad class of sampling designs within a unified metric framework.
We now introduce the corresponding concepts and definitions from a mathematical perspective. Let $\mathbf{k}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a bounded positive definite kernel with RKHS $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ and the associated canonical feature map $\Phi_{\mathbf{k}}(y):=\mathbf{k}(y,\cdot)\in\mathcal{H}$. For any $P\in\mathcal{M}_{b}(\mathcal{Y})$ in the space of finite signed Radon measures, define the kernel mean embedding as the mapping
@@
For our consistency arguments, we require, in addition, that $d_{\mathbf{k}}$ metrizes weak convergence on $\mathcal{M}_{b}(\mathcal{Y})$. That is, for any sequence $P_{1},\dots\in\mathcal{M}_{b}(\mathcal{Y})$, $P_{N}\rightsquigarrow P\iff d_{\mathbf{k}}(P_{N},\;P)\to 0$, as $N\to\infty$. We call kernels that satisfy this property universal (initially proposed in Steinwart2001-ja). This property concerns the approximation ability of the KME sample estimate to the true target as $N\to\infty$. We require the following assumptions
@@
    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of Sriperumbudur2010-xr).
@@
    $\mathbf{k}$ is $c_{0}$-universal. Meaning that the RKHS $\mathcal{H}$ induced by $\mathbf{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t. the uniform norm (Definition 2, Theorem 1 in Carmeli2010-ps).
@@
Assumptions (S1)–(K2) ensure that $\mu_{\mathbf{k}}(P)$ is well-defined for $P\in\mathcal{M}_{b}(\mathcal{Y})$ and that $d_{\mathbf{k}}$ is both discriminative (characteristic) and topologically compatible with weak convergence (metrizing). Assumption (K3) is not needed for the statistical theory, but when satisfied, admits a simple finite-dimensional approximation of the MMD through Fourier-feature representation for fast computation. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ Gretton_2007; More details are provided in Section S10, Supplementary Material.
Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding.
@@ 3 SDRF Algorithm and Theoretical Result
Random forests are machine learning methods that combine a CART-style donoho1997cart tree ensemble with a bagging step breiman1996bagging, and can also be viewed as kernel-type estimators (see, e.g. lin2006random; qiu2024random). Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems breiman2001random, and several works have documented their strong empirical performance diaz2006gene; fernandez2014we. More recently, conditional distributional versions of random forests based on kernel mean embeddings have been proposed for i.i.d. data cevid_michel_meinshausen_bühlmann_2022. Our proposed method, SDRF, constructs a kernel estimator $\widehat{\omega}_{\mathbf{x}}(\cdot)$ that directly targets conditional distributions and appropriately incorporates the survey design. Given the estimated neighborhood and weights (which together define $\widehat{\omega}_{\mathbf{x}}(\cdot)$), we can directly construct an estimator of the conditional distribution $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$, and hence a plug-in estimator for $\Psi(\mathbf{x})$, namely $\Psi(\mathbf{x})$: $\Psi(\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x}))$.
@@
4:  Draw $\forall k(i)$, $\mathbbm{1}(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
@@
11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
@@ 3.1 Bagging and Honesty
Bagging is a key ingredient of random forests breiman1996bagging, but its study is primarily conducted in an i.i.d. setting. In the i.i.d. case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators breiman1996bagging; Buhlmann2002-vp.
Under a survey design, however, Efron’s classical i.i.d. bootstrap Efron1993-eg can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
We adopt a design–aware bootstrap (BS) method in line with wang2022bootstrap as the default resampling method for SDRF. It generates multipliers $n^{*}$ that determine the additional up– or down–weighting applied to the sampled distribution, and satisfy properties (R1)–(R3) as detailed in Lemma 3.1.
@@
For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section S5.1 of the Supplement in the spirit of wang2022bootstrap (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
On resample multipliers $n^{*}$
@@
    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains the known cluster and strata structure. Each stratum is re–sampled separately, and if the design has two stages, the second stage resampling multipliers are independent across distinct PSUs conditional on first–stage resamples.
@@
The method by wang2022bootstrap belongs to the class of pseudo-population-based resampling methods. To summarize: For each re-sample, we first use design information such as $\{w_{i}\}_{i\in[n_{s}]}$ to re-construct via the multinomial distribution a pseudo finite population, with expected size equal to $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$; then we re-apply the sampling design $\mathbf{p}$ on the re-constructed finite population to obtain re-samples, thereby respecting the design structure. We refer interested readers to Mashreghi2016-nh; Conti_2022 for a review of common resampling methods under design.
For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ is used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ Athey2016-hs; Wager2018-lu; Athey2019-xy. With i.i.d. data, any disjoint partition of the observations yields honesty by construction. However, data from design carries possible dependency across subjects, so separation must be made at the PSU level to ensure the maximum level of independence conditioning on $\xi$. This motivates the introduction of $q\in(0,\,1)$. Alternatively, if one has access to data obtained from a continuous survey under the same design and is willing to assume that $P_{Y|X}$ stays unchanged between two consecutive time points, it is also possible to directly use one sample as $D_{est}$ and the other as $D_{split}$.
In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, Kish1995-aq) decreases as depth increases. As we discuss later, the performance of the MMD–based split criterion hinges on the effective sample size of $D_{split}$. In this case, one might consider spending the PSUs gradually over the tree growing process. See Algorithm 2, Supplementary Material for details.
@@ 3.2 Splitting Rule
The splitting criterion of an RF algorithm guides the formation of tree structure by determining how the feature space should be partitioned into increasingly homogeneous child nodes. Consider any node where a split is possible (has not met the stopping criteria). Denote $P_{a}\subseteq\mathcal{X}$ the feature space of that node. A candidate split can be parametrized as $\theta=(j,\;t)\in\Pi$, $j=1,\dots,p$, $t\in Range(X_{j})\subseteq[a_{j},\;b_{j}]$. This partitions $P_{a}$ into child regions $C_{L}(\theta):=\{x\in P_{a}:x_{j}\leq t\}$ and $C_{R}(\theta):=\{x\in P_{a}:\;x_{j}>t\}$. With RF being considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via a Hájek–type estimator. We find the split by maximizing the MMD between the KMEs of the distribution estimates between the two child regions. Formally, $\hat{\theta}=\arg\max_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)$, where
$M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},$
with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}\mathbbm{1}(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$. When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if they exist. We first provide a theorem justifying the choice of this split criterion.
@@
For a parent region $P_{a}\subseteq\mathcal{X}$ at depth $d$ generated by the tree-growing process, consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which define child regions $C_{L}(\theta)$ and $C_{R}(\theta)$. Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score, and let $M^{N}(\theta)$ be its finite-population analog. Assume (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3. Suppose further that $\mathcal{X}\subseteq\mathbb{R}^{p}$ is bounded. Then, conditional on the realized finite population,
@@
If, in addition, the finite population is an i.i.d. realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}\in\Pi$, $M(\theta)$ has a unique maximizer $\theta^{\dagger}\in\Pi$, and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta^{\dagger}$ under the super-population model).
@@
Theorem 3.2 provides a local result: for a fixed parent node $P_{a}$, the empirical MMD criterion $M^{*}_{n_{s}}(\theta)$ uniformly approximates its finite-population oracle counterpart $M^{N}(\theta)$ and also the super-population score $M(\theta)$. The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique van-der-Vaart2023-ur. The condition $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$ makes explicit that random feature subsampling (via mtrymtry) does not distort the split choice asymptotically, provided the candidate set becomes sufficiently dense around $\theta_{0}$ in expectation. Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to those obtained from using the full finite population, which is in particular assumption (B2) of Theorem 3.4 which is required for consistency in the MMD metric.
This raises the question that, since resamples are determined by non–negative multipliers $n^{*}$, can one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample? This could help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees. Proposition 3.3 below provides an affirmative answer to this question.